can someone help me with this subsutuion method problem please
Solve this system by elimination:
6x + 2y = 52
3x = 14 + 5y
x of solution (nearest integer):
y of solution (nearest integer):

Answer by Cromlix(4381)

6x + 2y = 52...1
3x - 5y = 14...2
Multiply (2) by 2
6x + 2y = 52...1
6x - 10y = 28...3
Subtract (3) from (2)
12y = 24
y = 2
Substitute y = 2
in Equation (1)
6x + 2y = 52....1
6x + 2(2) = 52
6x + 4 = 52
6x = 52 -4
6x = 48
x = 8
x = 8, y = 2
Hope this helps.
:-)